The graph of the equation $y =ax^2 + bx + c$, where $a$, $b$, and $c$ are constants, is a parabola with axis of symmetry $x = -3.$ Find $\frac{b}{a}$.

Answer by jim_thompson5910(35256)


x = h is the axis of symmetry
h = -b/(2a)
h = -3
So -3 = -b/(2a)

Multiply both sides by -2 and we get 6 = b/a which flips to b/a = 6

Answer: 6